OpenStudy (anonymous):

What are the zeros of f(x)=x^3+4x^2-7x-10

OpenStudy (cwrw238):

look like x-2 might be a factor (10 = 2*5) f(2) = 8 + 16 - 14 - 10 = 0 yup so divide f(x) by ( x -20 using long or synthetic division

OpenStudy (cwrw238):

* (x-2)

OpenStudy (cwrw238):

this will leave you with a trinomial to solve

OpenStudy (cwrw238):

can you do the division?